Two cards are draw at random from a well shuffled pack of 52cards. Find the probability that:
• Both are the same color
• Both are the different color
1
Expert's answer
2020-12-29T17:01:41-0500

The deck has cards of two colors: red and black. The probability to draw a black color is "\frac{26}{52}=0.5" and the same is the probability to draw a red color. The probability that two cards have a red color is: "\frac{26}{52}\cdot\frac{25}{51}\approx0.2451". The same is the probability that two cards have a black color. We used the multiplication rule (https://www.statisticshowto.com/multiplication-rule-probability/) The probability cards have the same color is: "0.2451\cdot2=0.4902". The probability that both cards have the different color is: "1-0.4902=0.5098"

Answer: the probability that two cards have the same color is: 0.4902; the probability that both cards have the different color is: 0.5098.